For Example, if John has a score of 72.5 on Exam-1 and 90 on Grade Frequenev Esan-2, he will have an increase the count of freauency valueA8 of Grade-A to 8 and Grade-C to 1. Q21: If we chose a student at random, what is the probability that student received grade-D? 1 Q22: If C or above is considered as a pass grade, and if Tota we chose a student at random, what is the probability that the student received a pass grade? Q23: What is the probability that the student selected at random got a fail grade? Q24: What is the probability that the student selected at random got a B or A grade? Q25: Is Q24 considered as mutually exclusive case? a) Yes b) No c) Not sure
Answer #1

21) Probability that student received grade D = 0

22) Number of students with C or above grades = 1 + 1 + 8 = 10

Thus, required probability to get a pass grade = 10/11

23) Probability of fail grade = 1/11

24) Probability of B or A grade = (1 + 8)/11 = 9/11

25) Yes, Q24 is mutually exclusive case as students can get a grade either in A or B